mohan covers 10.2km in 3hrs, the distance covered by him 5hrs
Answer by malaydassharma(59)

Mohan in 3 Hrs covers 10.2km

Mohan in 1 Hrs covers 10.2/3 km

Mohan in 5 Hrs covers 10.2 *5 /3 km = 3.4 * 5 = 17 Km